OpenStudy (anonymous):

name the property of 7(3+4)=21+28y

OpenStudy (anonymous):

you mean find y? 7(3+4)=21+28y 49 =21+28y 28 =28y y = 1

OpenStudy (anonymous):

I think the problem is supposed to be : name the property of 7(3 + 4y) = 21 + 28y. This is the distributive property.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."